OpenStudy (anonymous):

show that 1 and only 1 out of n, n+2, n+4 is divisible by 3 where n is any positive integer

OpenStudy (anonymous):

"show that 1 and only 1 out of n, n+2, n+4 " Could you draw that out? I don't know what you mean by that.

OpenStudy (anonymous):

The question means that for any value of n one of the functions above will be divisible by 3. For example if n = 1. \(n=1\) \(n+2=3\) \(n+4=5\) So for that example n+2 is divisible by 3. Does that make sense? Try to look at the difference between each function to get you started

OpenStudy (anonymous):

Start off with assuming there exists at least 2 out of them which are divisible by 3 and consider the difference between them, you would arrive at a contradiction.

OpenStudy (anonymous):

I was thinking about using mathematical induction.

mathslover (mathslover):

Using mathematical induction will be right @Azteck .. But there is an easy method too.

OpenStudy (anonymous):

Yeah I was thinking of the easy method, but I can't grasp it. I will see what you come up with mathslover.

mathslover (mathslover):

You have to wait for that... - 1 minute pls

OpenStudy (anonymous):

Yeah I will be on another tab helping someone else.

mathslover (mathslover):

Well by Euclid's division lemma we can write : \(n = 3q + r\) where q is quotient and r is remainder.

mathslover (mathslover):

We have \(0\le r < 3\)

mathslover (mathslover):

So we have 3 cases : r = 0 , r = 1 and r = 2 \(\textbf{Case 1:} \space n = 3q + 0\) ---- \(r = 0 \) \(n = 3q\) \(n+2 = 3q +2\)

mathslover (mathslover):

So it means when \(n+2\) is divided by 3 it gives remainder as 2 . Hence, n+2 is not divisible by 3 ------------ (1)

mathslover (mathslover):

Similarly : n + 4 = 3q + 4 that is : n +2 = 3q + 3 + 1 or : n + 4 = 3(q+1) + 1 or : n + 4 = 3m + 1 where m is an integer. thus n+4 is also NOT divisible by 3 .

mathslover (mathslover):

Hence when "n" is divisible by 3 , then it is NOT DIVISIBLE BY n+2 and n+4

mathslover (mathslover):

Similarly when n+2 is divisible by 3 then n and n+4 will not be divisible by 3. and when n+4 is divisible by 3 then n and n+2 will not be divisible by 3.

mathslover (mathslover):

Got it @queenbhavya , what are your thoughts @Azteck

OpenStudy (anonymous):

AH I see. Great work mathslover. Well thought out.

mathslover (mathslover):

:) @queenbhavya did you get what I did?

OpenStudy (anonymous):

thnkuuuuuu

OpenStudy (anonymous):

Actually considering the differences between any 2 of them is prolly the easiest way to do it, but if you are more familiar with this kind of proof then go for it.

mathslover (mathslover):

Best of Luck queenbhavya and : \[ \boxed{\boxed{\Large{\stackrel{\color{red}{\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star}}{\mathscr {\color{orange}{Welcome } \quad \color{green}{To}\quad \color{purple}{OpenStudy }!!}}}\\{\color{pink}{\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star}}}}\]

OpenStudy (unklerhaukus):

if n+4 is divisible by three , then so is n+1 n, n+1, n+2, are three consecutive integers so one of these will be divisible by three